Jon has 1121 pointsw in his math class. He must have 88% of the 1400 points possible by the end of the term to recieve credit for the class. What is the minimum number of additional points he must earn by the end of the term to receive credit for the class?
What I've tried:
88% of 1400=88/100*1400=1232-1121=111 points
The problem: 111+1121=1232, not 1400.
Where am I going wrong?

Answer by richwmiller(9135)

.88*1400=1232
1232-1121=111 more points to earn
111+1121=1232 which is 88% of 1400
1400 is 100%
He doesn't need 100% but only 88%
He needs 111 more points to have his 88%